If a 40-mL vial of sodium chloride is added to a 1-L container of water for injection, calculate the concentration of sodium chloride, in mEq/mL in the original vial, if the resultant dilution is 0.56% in strength.

PLEASE MAY HAVE A SOLUTION STEP BY STEP, I WILL APRECIATE
Answer #1

Initially 40 ml of sodium chloride (vial solution )contain= 5.85 g sodium chloride
Then 1 ml sodium chloride (vial solution) will contain = 5.85/40 g sodium chloride=0.146 g sodium chloride

The molecular weight of sodium chloride = 58.5
Than the equivalent weight of sodium chloride will be = 58.5mEq
The weight of sodium chloride = 58.5 ×10-3

= 0.0585 g

Given that the 0.0585 g of sodium chloride= 1 mEq sodium chloride
Then 0.146gm of sodium chloride = 1/0.0585 ×0.146= 2.495  mEq sodium chloride
= 2.5 mEq sodium chloride will be the strength